### v1

# Constraints on Compact Dark Matter Population from Micro-lensing Effect of Gravitational Wave for the third-generation gravitational Wave Detector

Abstract
Since the pioneering detection of gravitational wave (GW) from a binary black hole merger by the LIGO-Virgo collaboration, GW has become a powerful tool for probing astrophysics and cosmology. If compact dark matter (DM) candidates, e.g. primordial black holes, contribute a substantial fraction of the DM component across a broad mass spectrum, their gravitational influence would imprint distinctive micro-lensing signatures on long-wavelength GW signals. In this paper, we propose a systematic method ranging from simulating micro-lensing GWs to constraining population information of compact DM, combining future third-generation ground-based GW detector, i.e. Einstein Telescope, with a hierarchical Bayesian inference framework. This approach enables us to analyze micro-lensing signatures in GW events and probe the potential contribution of compact DM objects to the DM budget of universe. For a population with a power-law mass function, we demonstrate that detecting 8 micro-lensing GW events out of $10^{4}$ observed binary black holes would constrain the compact DM abundance to $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$.
preprint: APS/123-QED
I Introduction
The cosmic abundance of dark matter (DM), estimated to comprise $\sim 26\%$ of the total energy density in the universe, aligns with precision cosmological measurements (e.g. Cosmic Microwave Background anisotropies). Yet, despite this macroscopic confirmation, fundamental questions persist regarding its micro-physical properties and distribution at small scales. Particle physics and cosmology theories predict a variety of compact DM candidates, including massive compact halo objects (MACHOs), primordial black holes (PBHs), axion mini-clusters, compact mini-halos, boson stars, and fermion stars. In particular, PBH is taken as a potential candidate of DM and it has been a field of great astrophysical interest (see Sasaki et al. [1], Green and Kavanagh [2], Carr et al. [3] for recent reviews). To probe such compact objects, gravitational lensing effect is a powerful probe to constrain the abundance of compact DM over a broad mass range [4].
On September 14, 2015, the gravitational wave (GW) detectors Advanced LIGO [5, 6] and Virgo [7] observed the first GW signal from binary black hole (BBH), GW150914 [8]. In 2019, KAGRA’s baseline installation was finalized [9, 10, 11]. From then on, the LIGO, Virgo, and KAGRA network, which is referred to as the LVK Collaboration, detected approximately $\mathcal{O}(100)$ GW signals during the O1, O2, and O3 observing runs [12, 13, 14, 15]. As more and more GW events are detected, a revolutionary window into the universe was opened. GWs offer unprecedented opportunities to explore astrophysics and cosmology. Similar to electromagnetic waves, GWs can be lensed by massive object, such as galaxy, compact DM. However, unlike stellar light, GWs detected by ground-based observatories have much longer wavelengths, making wave optics crucial [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. This diffraction effect distorts the GW signal, creating waveform fringes. Since Jung and Shin [27] first proposed that the micro-lensing effect of GWs detected by advanced LIGO can probe the compact DM, the micro-lensing of GW has been extensively studied to constrain on the compact DM in follow-up researches [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. However, the most existing studies focus solely on constraining the upper limits of abundance of compact DM ($f_{\rm CO}\equiv\frac{\Omega_{\rm CO}}{\Omega_{\rm DM}}$) through null detection of micro-lensing GW observations [27, 28, 32, 30, 31, 33, 29, 34, 35]. In addition, the significant increase in the sensitivity of next-generation ground-based GW detectors, i.e. Einstein Telescope (ET) [40, 41] and Cosmic Explorer (CE) [42], will enable the detection of micro-lensing GW events, which would provide more valuable information into the properties of compact DM. In this work, we present a comprehensive framework, spanning from simulating micro-lensing GW signals to extracting population information of compact DM embedded in these signals. Furthermore, we provide quantitative projections for the detection capabilities of next-generation ground-based GW detectors i.e. ET.
[…]
II Methodology
In this Section, we present the theoretical framework for micro-lensing effects in GW, and develop a hierarchical Bayesian inference (HBI) approach to extract population properties of compact DM (e.g., mass function, and abundance) from an ensemble of micro-lensing GW events.
II.1 Theoretical framework for micro-lensing of gravitational wave
A GW is described by a small perturbation $h_{\mu\nu}$ over the background metric $g_{\mu\nu}^{(0)}$. The amplitude $h$ of this perturbation $h_{\mu\nu}$ satisfies the wave equation as
$\partial_{\mu}(\sqrt{-g^{(0)}}g^{(0)\mu\nu}\partial_{\nu}h)=0.$ (1)
Given micro-lensing events near the deflector, the lensed GW equation admits a Fourier-space approximation of the form
$(\nabla^{2}+4\pi^{2}f^{2})\bar{h}=16\pi^{2}f^{2}U\bar{h},$ (2)
where $U$ is the lensing potential in the background metric $g_{\mu\nu}^{(0)}$, and $\bar{h}$ is Fourier-transformed amplitude of $h$. We can define the dimensionless amplification factor as
$F(f)=\frac{\bar{h}_{\rm L}(f)}{\bar{h}_{0}(f)},$ (3)
where $\bar{h}_{\rm L}(f)$ and $\bar{h}_{0}(f)$ correspond to lensed and unlensed ($U=0$) signal, respectively. The solution of Eq. (3) in the thin lens approximation is given by
$F(f)=\frac{D_{\rm s}D_{\rm l}\theta_{\rm E}^{2}(1+z_{\rm l})f}{D_{\rm ls}i}\int d^{2}\vec{x}\leavevmode\nobreak\ \exp[2\pi ift_{\rm d}(\vec{x},\vec{y})],$ (4)
where $D_{\rm s}$, $D_{\rm l}$ and $D_{\rm ls}$ represent the angular diameter distance to the source, to the lens, and between the source and the lens, respectively. In addition, $\theta_{\rm E}$ is Einstein radius, and $\vec{y}$ is the impact position of source. $t_{\rm d}$ is the arrival time delay of the wave to the observer
$t_{d}(\vec{x},\vec{y})=\frac{D_{\rm s}D_{\rm l}\theta_{\rm E}^{2}(1+z_{\rm l})}{D_{\rm ls}}\bigg{[}\frac{1}{2}|\vec{x}-\vec{y}|^{2}-\psi(\vec{x})-\phi_{\rm m}(\vec{y})\bigg{]},$ (5)
where $\psi(\vec{x})$ is dimensionless deflection potential and $\phi_{\rm m}(\vec{y})$ is chosen such that the minimum arrival time delay is zero. For a point mass lens, we can use the rescaled position $x=|\vec{x}|$, $y=|\vec{y}|$ to get potential $\psi(\vec{x})=\ln x$ and $\phi_{m}(\vec{y})=(x_{m}-y)^{2}/2-\ln x_{m}$ with $x_{m}=(y+\sqrt{y^{2}+4})/2$. In this case, $F(f)$ can be written analytically as [19]
$\begin{split}F(f)=\exp\bigg{[}\frac{\pi\omega}{4}+i\frac{\omega}{2}\bigg{(}\ln\frac{\omega}{2}-2\phi_{\rm m}(y)\bigg{)}\bigg{]}\\ \times\Gamma\bigg{(}1-\frac{i\omega}{2}\bigg{)}{{}_{1}F_{1}}\bigg{(}i\frac{\omega}{2},1,\frac{i}{2}\omega y^{2}\bigg{)},\end{split}$ (6)
where $\omega=8\pi m_{\rm l}(1+z_{\rm l})f$ is a dimensionless parameter and $_{1}F_{1}$ is the confluent hypergeometric function. Obviously, the amplification factor of the point mass lens depends only on the redshifted mass of lens $M_{\rm l}^{z}\equiv(1+z_{\rm l})m_{\rm l}$ and the source position $y$.
II.2 Hierarchical Bayesian Inference
For population hyperparameters of compact DM $\Phi=[\bm{p}_{\rm mf},f_{\rm CO}]$ and $N_{\rm obs}$ detections of micro-lensing BBHs events $d=[d_{1},...,d_{N_{\rm obs}}]$, the likelihood follows an in-homogeneous Poisson distribution without any measurement uncertainty and selection effect is
$p(d|\Phi)\propto N(\Phi)^{N_{\rm obs}}e^{-N(\Phi)}\prod_{i}^{N_{\rm obs}}p_{\rm pop}(d_{i}|\Phi).$ (7)
However, with measurement uncertainty and selection effect taken into account, the likelihood for $N_{\rm obs}$ micro-lensing GW observations can be considered as [44, 45, 46]
$\begin{split}p(d|\Phi)\propto&N(\Phi)^{N_{\rm obs}}e^{-N_{\rm det}(\Phi)}\times\\ &\prod_{i}^{N_{\rm obs}}\int dm\,L(d_{i}|m_{\rm l})p_{\rm pop}(m_{\rm l}|\Phi),\end{split}$ (8)
where the likelihood of one micro-lensing event $L(d_{i}|m_{\rm l})$ is proportional to the posterior $p(m_{\rm l}|d_{i})$. $N(\Phi)$ is the total number of micro-lensing events in the model characterized by the set of population parameters $\Phi$ as
$\begin{split}N(\Phi)=&\int dm_{\rm l}\int dz_{\rm s}\int_{0}^{z_{\rm s}}dz_{\rm l}\frac{dn(m_{\rm l},\Phi)}{dm_{\rm l}}\times\\ &\frac{d\chi(z_{\rm l})}{dz_{\rm l}}(1+z_{\rm l})^{2}\sigma(m_{\rm l},z_{\rm l},z_{\rm s})N_{\rm s}P_{\rm s}(z_{\rm s}),\end{split}$ (9)
where $\chi(z_{\rm l})$ is the comoving distance, $N_{\rm s}$ and $P_{\rm s}(z_{\rm s})$ respectively represent the number and redshift distribution of detectable GW sources, ${dn(m_{\rm l},\Phi)}/{dm_{\rm l}}$ is the comoving number density of the compact DM in a certain extended mass distribution $\psi(m_{\rm l},\bm{p}_{\rm mf})$
$\frac{dn(m_{\rm l},\Phi)}{dm_{\rm l}}=\frac{f_{\rm CO}\Omega_{\rm DM}\rho_{\rm c}}{m_{\rm l}}\psi(m_{\rm l},\bm{p}_{\rm mf}),$ (10)
where $\Omega_{\rm DM}$ and $\rho_{\rm c}$ are the current fractional contribution of DM to the total energy density and critical density of the universe. In Eq. (9), $\sigma(m_{\rm l},z_{\rm l},z_{\rm s})$ is the lensing cross section
$\sigma(m_{\rm l},z_{\rm l},z_{\rm s})=\frac{4\pi m_{\rm l}D_{\rm l}D_{\rm ls}}{D_{\rm s}}y_{0}^{2}.$ (11)
It should be emphasized that we take the impact parameter as $y_{0}=5$, because $y_{0}>5$ are difficult to be identified as micro-lensing signals [29, 31]. In addition, $p_{\rm pop}(m_{\rm l}|\Phi)$ is the normalized distribution of lens masses and written as
$p_{\rm pop}(m_{\rm l}|\Phi)=\frac{1}{N(\Phi)}\frac{dN(m_{\rm l},\Phi)}{dm_{\rm l}}=\psi(m_{\rm l},\bm{p}_{\rm mf}).$ (12)
Meanwhile, $N_{\rm det}(\Phi)$ is the number of detectable micro-lensing GW events and can be defined as
$\begin{split}N_{\rm det}(\Phi)=&\int dm_{\rm l}\int dz_{\rm s}\int_{0}^{z_{\rm s}}dz_{\rm l}\frac{dn(m_{\rm l},\Phi)}{dm_{\rm l}}\times\\ &\frac{d\chi(z_{\rm l})}{dz_{\rm l}}(1+z_{\rm l})^{2}\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})N_{\rm s}P_{\rm s}(z_{\rm s}),\end{split}$ (13)
where $\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})$ is the cross section that micro-lensing signal can be detected and depends on Eq. (30) via
$\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})=\frac{4\pi m_{\rm l}D_{\rm l}D_{\rm ls}}{D_{\rm s}}y_{\max}(\lambda)^{2}.$ (14)
where $\lambda\in[M_{\rm l}^{z},z_{\rm s},m_{1},m_{2},$......] contains the parameter sets of the GW source and lens. Meanwhile, we can define the detected fraction $\xi(\Phi)$ of micro-lensing GWs as
$\xi(\Phi)\equiv N_{\rm det}(\Phi)/N(\Phi),$ (15)
Then the posterior distribution $p(\Phi|d)$ can be calculated from
$p(\Phi|d)=\frac{p(d|\Phi)p(\Phi)}{Z_{\mathcal{M}}},$ (16)
where $p(\Phi)$ is the prior distribution for population hyperparameters $\Phi$. In addition, $Z_{\mathcal{M}}$ is both the normalized factor and the Bayesian evidence for the population model $\mathcal{M}$. This normalized factor can be calculated from the integral of the numerator of Eq. (16) over $\Phi$, i.e.
$Z_{\mathcal{M}}=\int d\Phi\leavevmode\nobreak\ p(d|\Phi)p(\Phi).$ (17)
Figure 1: Left: The posterior distribution for one of lensing GW events in 8 detectable micro-lensing GW events. Right: The cross represents the range of values within $2\sigma$ confidence level of luminosity distance for GW sources and redshift mass for lens extracted from the posterior distribution of 8 detectable micro-lensing GW events. And the scattered points represent the true values of these micro-lensing events.
III Simulation
In this Section, we introduce the simulation of GWs and how to generate micro-lensing GW events.
III.1 Simulation for model of GW sources
Although understanding the origin of GWs of BBHs is still an open issue, we select BBH merger events as the GW sources for this investigation due to some reasons: These systems exhibit high signal-to-noise ratio (SNR), robust waveform templates, and high event rates. The redshift distribution of GW sources is drawn from the phenomenological fit $R_{\rm BBH}$ to the population synthesis rate
$P_{\rm BBH}(z_{\rm s})=\frac{1}{Z_{\rm BBH}}\frac{R_{\rm BBH}(z_{\rm s})}{1+z_{\rm s}}\frac{dV_{\rm c}}{dz_{\rm s}},$ (18)
where $Z_{\rm BBH}$ is the normalization constant, and $R_{\rm BBH}$ is the merger rate per unit of volume in the source frame and takes the form [30, 47]
$R_{\rm BBH}(z_{\rm s})=R_{0}\int_{t_{\min}}^{t_{\max}}R_{\rm f}(t(z_{\rm s})-t_{\rm d})P(t_{\rm d})dt_{\rm d},$ (19)
where $t_{\min}=50\leavevmode\nobreak\ {\rm Myr}$, $t_{\max}$ is set to the Hubble time, $P(t_{\rm d})\propto t_{\rm d}^{-1}$ is the time delay distribution which is set the power-law, and $R_{\rm f}(t(z_{\rm s})-t_{\rm d})$ is the Star Formation Rate (SFR) as [48, 30]
$R_{\rm f}(z_{\rm s})=[1+(1+z_{\rm p})^{-\lambda-\kappa}]\frac{(1+z_{\rm s})^{\lambda}}{1+\bigg{(}\frac{1+z_{\rm s}}{1+z_{\rm p}}\bigg{)}^{\lambda+\kappa}},$ (20)
where we set $\lambda=2.6$, $\kappa=3.6$, and $z_{\rm p}=2.2$. We use a power-law mass distribution for the heavier black hole in the BBHs as [28, 29, 30, 31]
$P_{\rm BBH}(m_{\rm s})=Z_{\rm m}m_{\rm s}^{-\alpha}\mathcal{H}(m_{\rm s}-m_{\min})\mathcal{H}(m_{\max}-m_{\rm s}),$ (21)
where $\mathcal{H}$ is the Heaviside step function and $Z_{\rm m}$ is the normalization constant as
$Z_{\rm m}=\left\{\begin{aligned} \frac{1-\alpha}{M_{\max}^{1-\alpha}-M_{\min}^{1-\alpha}},\quad\alpha\neq 1,\\ \frac{1}{\ln(M_{\max}/M_{\min})},\quad\alpha=1.\end{aligned}\right$. (22)
The heavier black hole $m_{1}$ satisfies the mass distribution as Eq.(21) with $\alpha=2.35$ in the range of $m_{1}\in[5,100]\leavevmode\nobreak\ M_{\odot}$, and the mass of $m_{2}$ uniformly distributes in the interval $m_{1}/18,m_{1}]$  [^1] For other parameters in GW sources, such as the orbital inclination and the merging orbital phase, we adopt a uniform distribution. In the following discussion, we do not show them, but we all take their influence into consideration..
III.2 Simulation for micro-lensing events
To ensure the physical rationality and computational efficiency of micro-lensing GW events simulation, we have developed a five-step process:
  • Generating the detectable BBH events: Based on the redshift distribution and mass distribution in above subsection, we can simulate $10^{4}$ detectable unlensed BBH events with $\rm{SNR}\geq 8$ for ET. The SNR of the unlensed GW signal can be computed as
    ${\rm{SNR}_{0}}=\left\langle\bar{h}_{0}|\bar{h}_{0}\right\rangle=\sqrt{4\int\frac{|\bar{h}_{0}(f)|^{2}}{S_{\rm n}(f)}df},$ (23)
    where $S_{\rm n}(f)$ is the power spectral density for ET.
  • Generating the micro-lensing GW events: By assuming that the compact DM is distributed uniformly in comving volume, the probability that a GW source at $z_{\rm s}$ is lensed by an intervening compact DM object can be given by optical depth as
    $P_{\rm L}(z_{\rm s})=1-\exp(-\tau(z_{\rm s})).$ (24)
    We identify a BBH GW signal as a lensed event when $P_{\rm L}(z_{\rm s})$ is larger than a random number uniformly distribution between 0 and 1. Therein the optical depth $\tau(z_{\rm s})$ can be written as
    $\begin{split}\tau(z_{\rm s})=&\int_{0}^{+\infty}dm_{\rm l}\int_{0}^{z_{\rm s}}dz_{\rm l}\int_{0}^{y_{0}}dy\\ &\tau(m_{\rm l},z_{\rm l},y),\end{split}$ (25)
    where $y_{0}$ is the maximum impact parameter. $\tau(m_{\rm l},z_{\rm l},y)$ is the differential optical depth for the mass of lens $m_{\rm l}$, redshift of lens $z_{\rm l}$ and impact parameter $y$, given by
    $\begin{split}\tau(m_{\rm l},z_{\rm l},y)=&3f_{\rm CO}\Omega_{\rm DM}y\psi(m_{\rm l},\bm{p}_{\rm mf})\times\\ &\frac{H_{0}^{2}(1+z_{\rm l})^{2}}{H(z_{\rm l})}\frac{D_{\rm l}D_{\rm ls}}{D_{\rm s}},\end{split}$ (26)
    where we set $f_{\rm CO}=10^{-2.5}$.
  • Generating the parameters of lens: When a GW event is identified as lensed one, the value of $z_{\rm l}$ is adopted from a probability distribution $P(z_{\rm l})$ given by Eq. (26)
    $P(z_{\rm l})=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{z_{\rm s}}dz_{\rm l}\leavevmode\nobreak\ \tau(m_{\rm l},z_{\rm l},y)}.$ (27)
    In addition, the distribution $P(y)$ also can be obtained from Eq. (26) in the range of $y\in[0,y_{0}]$
    $P(y)=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{y_{0}}dy\leavevmode\nobreak\ \tau(m_{\rm l},z_{\rm l},y)}=\frac{2}{y_{0}^{2}}y.$ (28)
    We set $y_{0}=5$ as mentioned above. For extended mass distribution of lens mass $\psi(m_{\rm l},\bm{p}_{\rm mf})$ same as in Eq. (10), we assume the maximum value of the extended power-law mass function is $M_{\max}=10^{3}\leavevmode\nobreak\ M_{\odot}$ and the minimum value is $M_{\min}=10\leavevmode\nobreak\ M_{\odot}$. The mass of the lens satisfies the power-law distribution described as
    $\begin{split}\psi(m_{\rm l},\gamma)=&\mathcal{N}_{\rm pl}m_{\rm l}^{\gamma-1}\mathcal{H}(m_{\rm l}-M_{\min})\times\\ &\mathcal{H}(M_{\max}-m_{\rm l}),\end{split}$ (29)
    where $\mathcal{N}_{\rm pl}$ is the normalization, and we set the power-law index $\gamma=1$.
  • Computing lensed criterion: After generating unlensed and lensed GW waveform with simulated parameters for sources and lens, we use the mismatch $\epsilon$ and $\rm{SNR_{L}}$ as criterion for selection of lensing GW following [25]
    $\epsilon\times{\rm{SNR_{L}}}^{2}>1,$ (30)
    where mismatch $\epsilon$ is expressed as
    $\epsilon=1-\frac{\left\langle\bar{h}_{\rm L}|\bar{h}_{0}\right\rangle}{\sqrt{\left\langle\bar{h}_{\rm L}|\bar{h}_{\rm L}\right\rangle\left\langle\bar{h}_{0}|\bar{h}_{0}\right\rangle}}.$ (31)
  • Extract parameters of micro-lensing GWs: We use the Bayes’ theorem to forecast the constraints of the parameters $m_{1},m_{2},d_{\rm L}(z_{\rm s}),M_{\rm l}^{z},y]$ from lensed GW signals. The likelihood for each GW signals can be written in the form of an inner product [39]
    $p(d|\theta)\propto\exp\bigg{[}-\frac{1}{2}\left\langle d-\bar{h}(\theta)|d-\bar{h}(\theta)\right\rangle^{2}\bigg{]},$ (32)
    where $\bar{h}(\theta)$ represents the model of micro-lensing GW waveform.
Figure 2: Left: The proportion of detectable micro-lensing GW events for Eq. (15). The blue data points represent simulated values, while the solid red line indicates the corresponding polynomial fit to these data. Right: Solid lines with points represent inferred lens masses of 8 micro-lensing GWs at $95\%$ confidence levels with median value derived from the posterior distribution of redshifted lens mass $p(M_{\rm l}^{z}|d_{i})$ and luminosity-distance distribution of GW sources $p(d_{\rm L}(z_{\rm s})|d_{i})$ in Figure 1. The star points represent the true values of the simulated lens masses.
From $10^{4}$ simulated GW events of BBHs, we identify 8 detectable micro-lensing GW signals. Moreover, we present the posterior distribution of the parameters of these events in Figure 1. As shown in the right panel of Figure 1, we compare the true simulated values of the luminosity distance for GW source and the redshift mass of lens with their posterior-derived counterparts. These results demonstrate good agreement within $2\sigma$ confidence levels.
IV Results
The selection effects of gravitational lensing systems and the uncertainties in observed lens masses are presented in Figure 2. Specifically, the left panel of Figure 2 shows the detectable fraction of lensed GW events as a function of the power-law index. Our analysis reveals a positive correlation between the power-law index and the detection rate of micro-lensing GW events. This trend is physically expected, as larger $\gamma$ values correspond to an increased abundance of higher-mass lenses, which generally produce stronger lensing signals with higher SNRs. In addition, as shown in the right panel of Figure 2, we present the inferred lens masses and their true values of the simulated lens. In analysis for extract parameters of micro-lensing GWs, the posterior distributions of both the luminosity distance of the source $p(d_{\rm L}(z_{\rm s})|d_{i})$ and the lens redshifted mass $p(M_{\rm l}^{z}|d_{i})$ can be obtained for each event. By combining these posterior distributions with optical depth theory in Eq. (27), we can derive the likelihood function of the lens mass $L(d_{i}|m_{\rm l})$ for each lensing GW events. Our analysis demonstrates that the inferred lens masses are consistent with their true values within $2\sigma$ confidence levels.
Figure 3: The posterior distributions for hyperparameters $\gamma,\log_{10}(f_{\rm CO})]$ within $2\sigma$ confidence level. The blue solid point represent the truth values for hyperparameters $\gamma=1,\log_{10}(f_{\rm CO})=-2.5]$ in our simulation.
For the population hyperparameters of compact DM, we fix the minimum mass $M_{\min}=10\leavevmode\nobreak\ M_{\odot}$ and maximum mass $M_{\max}=10^{3}\leavevmode\nobreak\ M_{\odot}$ in the mass function of compact DM, and we adopt uniform priors for the power-law index $\gamma\in\mathcal{U}[-1,3]$ and abundance of compact DM $f_{\rm CO}\in\mathcal{U}[-4,-1]$ in Eq. (16), respectively. Then we incorporate the above-mentioned 8 inferred lens masses from micro-lensing GW events into the EMCEE [49] with the posterior Eq. (16) to estimate population hyperparameters $\Phi\in[\gamma,\leavevmode\nobreak\ f_{\rm CO}]$. Our results are shown in Figure. 3. We obtain that the best-fit values and $68\%$ confidence levels for the hyperparameters $\gamma,\leavevmode\nobreak\ f_{\rm CO}]$ are $\gamma=0.41^{+0.42}_{-0.37}$, $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$, respectively. As shown by the blue data point in Figure 3, which indicate the truth values for hyperparameter $\Phi\in[\gamma=1,\leavevmode\nobreak\ \log_{10}(f_{\rm CO})=-2.5]$ in our simulations, our posterior estimates demonstrate consistency with these input truth values within $2\sigma$ confidence levels.
V Conclusion
In this paper, we present a comprehensive methodology for: (1) simulating GW events mciro-lensed by compact DM, (2) extracting information about both the GW sources and lens for each micro-lensing events, and (3) constraining the population properties of compact DM from these events. Specifically, by combining the future third-generation ground-based GW detector ET with point mass model, we calculate the posterior distribution of GW sources and lens parameters for the detection of 8 micro-lensing GW events among $10^{4}$ observed BBH mergers. Then, based on the hierarchical Bayesian inference framework, we demonstrate that for a population characterized by a power-law mass function, these 8 micro-lensing GW events would constrain the power-law index to $\gamma=0.41^{+0.42}_{-0.37}$ and compact DM abundance to $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$ at $68\%$ confidence levels as shown in Figure 3. Our analysis demonstrates agreement between the derived results and the true input values from simulations within $2\sigma$ confidence levels. This discrepancy is expected, as our theoretical predictions suggest that more than 10 micro-lensing GW events should be detectable given the true parameter values. However, only 8 micro-lensing GW events were detected, which explains why the derived constraints on both $f_{\rm CO}$ and $\gamma$ are lower than the true input values. The implementation of our methodology will play a pivotal role in future observations of micro-lensing GWs: it enables not only the precise extraction of population characteristics of compact DM, but also provides a pathway to probe their fundamental nature—whether they stem from primordial black holes, exotic compact objects, or other non-particle dark matter candidates.
The constraints on compact DM will improve significantly due to the rapidly growing number of micro-lensing GW detections from upcoming surveys and the enhanced horizon distance of next-generation GW detectors. This advancement will lead to substantial overlap between the parameter spaces probed by GW observations and those constrained by electromagnetic observations. Consequently, combining these two powerful multi-messenger approaches will enable joint constraints on the abundance and mass distribution of compact DM. Such joint analyses are expected to play a pivotal role in unraveling the nature of compact DM and their formation mechanisms, shedding light on physics in the early universe.
VI Acknowledgements
This work is supported by National Key R&D Program of China under Grant No.2024YFC2207400; China National Postdoctoral Program for Innovative Talents under Grant No.BX20230271; National Key R&D Program of China Grant No. 2021YFC2203001; National Natural Science Foundation of China under Grants Nos.12447137, 11920101003, 12021003, 11633001, 12322301, and 12275021; the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant Nos. XDB2300000 and the Interdiscipline Research Funds of Beijing Normal University.

### v2

@@
 Abstract
-Since the pioneering detection of gravitational wave (GW) from a binary black hole merger by the LIGO-Virgo collaboration, GW has become a powerful tool for probing astrophysics and cosmology.
+Since the pioneering detection of gravitational wave (GW) from a binary black hole merger by the LIGO-Virgo collaboration, GW has become a powerful probe for astrophysics and cosmology.
 If compact dark matter (DM) candidates, e.g.
-primordial black holes, contribute a substantial fraction of the DM component across a broad mass spectrum, their gravitational influence would imprint distinctive micro-lensing signatures on long-wavelength GW signals.
-In this paper, we propose a systematic method ranging from simulating micro-lensing GWs to constraining population information of compact DM, combining future third-generation ground-based GW detector, i.e.
-Einstein Telescope, with a hierarchical Bayesian inference framework.
-This approach enables us to analyze micro-lensing signatures in GW events and probe the potential contribution of compact DM objects to the DM budget of universe.
-For a population with a power-law mass function, we demonstrate that detecting 8 micro-lensing GW events out of $10^{4}$ observed binary black holes would constrain the compact DM abundance to $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$.
+primordial black holes, contribute a substantial fraction of the DM component across a broad mass range, they would yield distinctive micro-lensing signatures on GW signals.
+In this paper, based on the third-generation ground-based GW detector, i.e.
+Einstein Telescope, we propose to constrain population information of compact DM by simulating micro-lensing GWs and analyzing with the hierarchical Bayesian inference framework.
+For a population with a power-law mass function, we demonstrate that detections of several micro-lensing GW signals in $10^{4}$ binary black holes coalescence events would constrain the abundance of compact DM to $\sim 10^{-3}$.
+It suggests that searching for and identifying micro-lensing signatures in future detections could be complementary and helpful in constraining compact DM scenarios.
 preprint: APS/123-QED
 I Introduction
-The cosmic abundance of dark matter (DM), estimated to comprise $\sim 26\%$ of the total energy density in the universe, aligns with precision cosmological measurements (e.g.
-Cosmic Microwave Background anisotropies).
-Yet, despite this macroscopic confirmation, fundamental questions persist regarding its micro-physical properties and distribution at small scales.
-Particle physics and cosmology theories predict a variety of compact DM candidates, including massive compact halo objects (MACHOs), primordial black holes (PBHs), axion mini-clusters, compact mini-halos, boson stars, and fermion stars.
+The cosmic abundance of dark matter (DM) has been estimated by precision cosmological measurements, such as the analysis of Cosmic Microwave Background anisotropies, to be approximately $26\%$ of the total energy density of the universe [1].
+Despite this macroscopic confirmation, fundamental questions persist regarding its microphysical properties and distribution at small scales.
+Various hypotheses for dark matter may imply the presence of a variety of compact DM candidates, including primordial black holes (PBHs) [2, 3, 4], mini-halos [5, 6], boson stars [7, 8], and fermion stars [9].
 In particular, PBH is taken as a potential candidate of DM and it has been a field of great astrophysical interest (see Sasaki et al.
@@
 GWs offer unprecedented opportunities to explore astrophysics and cosmology.
-Similar to electromagnetic waves, GWs can be lensed by massive object, such as galaxy, compact DM.
-However, unlike stellar light, GWs detected by ground-based observatories have much longer wavelengths, making wave optics crucial [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
+Similar to electromagnetic waves, GWs can be lensed by massive objects, such as galaxies, compact DM.
+However, unlike stellar light, GWs detected by observatories have much longer wavelengths, making wave optics crucial [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
 This diffraction effect distorts the GW signal, creating waveform fringes.
 Since Jung and Shin [39] first proposed that the micro-lensing effect of GWs detected by Advanced LIGO can probe the compact DM, the micro-lensing of GW has been extensively studied to constrain on the compact DM in follow-up researches [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51].
-However, the most existing studies focus solely on constraining the upper limits of abundance of compact DM ($f_{\rm CO}\equiv\frac{\Omega_{\rm CO}}{\Omega_{\rm DM}}$) through null detection of micro-lensing GW observations [27, 28, 32, 30, 31, 33, 29, 34, 35].
+However, the most existing studies focus solely on constraining the upper limits of abundance of compact DM ($f_{\rm CO}\equiv\Omega_{\rm CO}/\Omega_{\rm DM}$) through null detection of micro-lensing GW observations [39, 40, 44, 42, 43, 45, 41, 46, 47].
 In addition, the significant increase in the sensitivity of next-generation ground-based GW detectors, i.e.
-Einstein Telescope (ET) [40, 41] and Cosmic Explorer (CE) [42], will enable the detection of micro-lensing GW events, which would provide more valuable information into the properties of compact DM.
+Einstein Telescope (ET) [52, 53] and Cosmic Explorer (CE) [54], will enable the detection of micro-lensing GW events, which would provide more valuable information about the properties of compact DM.
 In this work, we present a comprehensive framework, spanning from simulating micro-lensing GW signals to extracting population information of compact DM embedded in these signals.
@@
 […]
-II Methodology
-In this Section, we present the theoretical framework for micro-lensing effects in GW, and develop a hierarchical Bayesian inference (HBI) approach to extract population properties of compact DM (e.g., mass function, and abundance) from an ensemble of micro-lensing GW events.
-II.1 Theoretical framework for micro-lensing of gravitational wave
-A GW is described by a small perturbation $h_{\mu\nu}$ over the background metric $g_{\mu\nu}^{(0)}$. The amplitude $h$ of this perturbation $h_{\mu\nu}$ satisfies the wave equation as
+II Theoretical framework for micro-lensing of gravitational wave
+To describe the lensing effect of GWs, we consider gravitational potential of the lens $U(\vec{r})$ in the FLRW background $g_{\mu\nu}^{(0)}$,
+$ds^{2}=-(1+2U(\vec{r}))dt^{2}+a(t)^{2}(1-2U(\vec{r}))d\vec{r}^{2}.$ (1)
+GW is described by a small perturbation $h_{\mu\nu}$ over the background metric $g_{\mu\nu}^{(0)}$. The amplitude $h$ of this perturbation $h_{\mu\nu}$ satisfies the wave equation as
 $\partial_{\mu}(\sqrt{-g^{(0)}}g^{(0)\mu\nu}\partial_{\nu}h)=0.$ (2)
-Given micro-lensing events near the deflector, the lensed GW equation admits a Fourier-space approximation of the form
+As a result of lensing being approximable as a local event occurring near the lens object, we can neglect the time dependence of the scale factor and write the Helmholtz equation for $h$ in Fourier space as
 $(\nabla^{2}+4\pi^{2}f^{2})\bar{h}=16\pi^{2}f^{2}U\bar{h},$ (3)
-where $U$ is the lensing potential in the background metric $g_{\mu\nu}^{(0)}$, and $\bar{h}$ is Fourier-transformed amplitude of $h$. We can define the dimensionless amplification factor as
+where $\bar{h}$ is Fourier-transformed amplitude of $h$. We can define the dimensionless amplification factor as
 $F(f)=\frac{\bar{h}_{\rm L}(f)}{\bar{h}_{0}(f)},$ (4)
-where $\bar{h}_{\rm L}(f)$ and $\bar{h}_{0}(f)$ correspond to lensed and unlensed ($U=0$) signal, respectively. The solution of Eq. (3) in the thin lens approximation is given by
-$F(f)=\frac{D_{\rm s}D_{\rm l}\theta_{\rm E}^{2}(1+z_{\rm l})f}{D_{\rm ls}i}\int d^{2}\vec{x}\leavevmode\nobreak\ \exp[2\pi ift_{\rm d}(\vec{x},\vec{y})],$ (4)
-where $D_{\rm s}$, $D_{\rm l}$ and $D_{\rm ls}$ represent the angular diameter distance to the source, to the lens, and between the source and the lens, respectively.
-In addition, $\theta_{\rm E}$ is Einstein radius, and $\vec{y}$ is the impact position of source.
-$t_{\rm d}$ is the arrival time delay of the wave to the observer
-$t_{d}(\vec{x},\vec{y})=\frac{D_{\rm s}D_{\rm l}\theta_{\rm E}^{2}(1+z_{\rm l})}{D_{\rm ls}}\bigg{[}\frac{1}{2}|\vec{x}-\vec{y}|^{2}-\psi(\vec{x})-\phi_{\rm m}(\vec{y})\bigg{]},$ (5)
+where $\bar{h}_{\rm L}(f)$ and $\bar{h}_{0}(f)$ correspond to lensed and unlensed ($U=0$) signal, respectively.
+The solution of Eq.
+(4) in the thin-lens and point source approximation is given by [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]
+$F(f)=\frac{D_{\rm s}\xi_{0}^{2}(1+z_{\rm l})f}{iD_{\rm ls}D_{\rm l}}\int d^{2}\vec{x}\penalty 10000\ \exp[2\pi ift_{\rm d}(\vec{x},\vec{y})],$ (5)
+where $\xi_{0}\equiv\theta_{\rm E}D_{\rm l}$ is a normalization factor with Einstein radius $\theta_{\rm E}$ to match the characteristic scale of the lensing configuration, $D_{\rm l}$ and $D_{\rm ls}$ represent the angular diameter distance to the source, to the lens, and between the source and the lens, respectively.
+In addition, the dimensionless vectors $\vec{x}$ and $\vec{y}$ are defined as
+$\begin{split}&\vec{x}\equiv\frac{\vec{\xi}}{\xi_{0}},\\ &\vec{y}\equiv\frac{D_{\rm l}}{D_{\rm s}\xi_{0}}\vec{\eta},\end{split}$ (6)
+where $\vec{\xi}$ and $\vec{\eta}$ are the physical coordinates of the image in the lens plane and of the source in the source plane, respectively.
+$\Delta t_{\rm d}$ is the arrival time delay of the wave to the observer
+$\begin{split}&\Delta t_{d}(\vec{x},\vec{y})=\frac{D_{\rm s}\xi_{0}^{2}(1+z_{\rm l})}{D_{\rm l}D_{\rm ls}}\times\\ &\bigg[\frac{1}{2}|\vec{x}-\vec{y}|^{2}-\psi(\vec{x})-\phi_{\rm m}(\vec{y})\bigg],\end{split}$ (7)
 where $\psi(\vec{x})$ is dimensionless deflection potential and $\phi_{\rm m}(\vec{y})$ is chosen such that the minimum arrival time delay is zero.
@@
 where $\omega=8\pi m_{\rm l}(1+z_{\rm l})f$ is a dimensionless parameter and $_{1}F_{1}$ is the confluent hypergeometric function.
-Obviously, the amplification factor of the point mass lens depends only on the redshifted mass of lens $M_{\rm l}^{z}\equiv(1+z_{\rm l})m_{\rm l}$ and the source position $y$.
-II.2 Hierarchical Bayesian Inference
-For population hyperparameters of compact DM $\Phi=[\bm{p}_{\rm mf},f_{\rm CO}]$ and $N_{\rm obs}$ detections of micro-lensing BBHs events $d=[d_{1},...,d_{N_{\rm obs}}]$, the likelihood follows an in-homogeneous Poisson distribution without any measurement uncertainty and selection effect is
+Obviously, the amplification factor of the point mass lens only depends on the redshifted mass of lens $M_{\rm l}^{z}\equiv(1+z_{\rm l})m_{\rm l}$ and the source position $y$.
+The geometrical optics limit corresponds to the scenario where the GW wavelength is significantly smaller than the Schwarzschild radius of the lens.
+In this regime, the amplification factor is well approximated by a sum of discrete magnification contributions from multiple images as
+$\begin{split}F(f)=|\mu_{+}|^{1/2}-i|\mu_{-}|^{1/2}\exp(i2\pi f\Delta t_{\rm d}),\end{split}$ (9)
+where $\mu_{+}$ and $\mu_{-}$ are the magnifications of the two lensing images.
+Figure 1: Left: The posterior distribution for one of lensing GW events in 8 detectable micro-lensing GW events.
+Right: The cross represents the range of values within $2\sigma$ confidence level of luminosity distance for GW sources and redshift mass for lens extracted from the posterior distribution of 8 detectable micro-lensing GW events.
+And the scattered points represent the true values of these micro-lensing events.
+III Simulation
+In this section, we introduce the simulation of GWs and how we generate micro-lensing GW events.
+To ensure the physical rationality and computational efficiency of micro-lensing GW events simulation, we have developed a five-step process:
+  • Generating the detectable BBH events: Although understanding the astrophysical origin of GWs of BBHs is still an open issue, we select BBH merger events as the GW sources for this investigation due to these reasons: These systems exhibit high signal-to-noise ratio (SNR), robust waveform templates, and high event rates.
+In addition, most GW sources observed so far by LVK Collaboration came from BBH mergers.
+The redshift distribution of GW sources is drawn from the phenomenological fit $R_{\rm BBH}$ to the population synthesis rate
+    $P_{\rm BBH}(z_{\rm s})=\frac{1}{Z_{\rm BBH}}\frac{R_{\rm BBH}(z_{\rm s})}{1+z_{\rm s}}\frac{dV_{\rm c}}{dz_{\rm s}},$ (10)
+    where $Z_{\rm BBH}$ is the normalization constant, and $R_{\rm BBH}$ is the merger rate per unit of volume in the source frame and takes the form [42, 55]
+    $\begin{split}&R_{\rm BBH}(z_{\rm s})=R_{0}\times\\ &\int_{t_{\min}}^{t_{\max}}R_{\rm f}(t(z_{\rm s})-t_{\rm d})P(t_{\rm d})dt_{\rm d},\end{split}$ (11)
+    where $t_{\min}=50\penalty 10000\ {\rm Myr}$, $t_{\max}$ is set to the Hubble time, $P(t_{\rm d})\propto t_{\rm d}^{-1}$ is the time delay distribution which is set the power-law, and $R_{\rm f}(t(z_{\rm s})-t_{\rm d})$ inherits the redshift dependence of Star Formation Rate (SFR) as [56, 42]
+    $\begin{split}R_{\rm f}(z_{\rm s})=[1+(1+z_{\rm p})^{-\lambda-\kappa}]\times\\ \frac{(1+z_{\rm s})^{\lambda}}{1+\bigg(\frac{1+z_{\rm s}}{1+z_{\rm p}}\bigg)^{\lambda+\kappa}},\end{split}$ (12)
+    if $z_{\rm p}$ is much larger than 0 and $\lambda+\kappa$ is positive, this model is expected to behave like $(1+z)^{\lambda}$ and $(1+z)^{-\kappa}$ at low and high redshifts respectively.
+In Eq.
+(12), we set $\lambda=2.6$, $\kappa=3.6$, and $z_{\rm p}=2.2$ [56].
+For mass function of GW sources, we use a power-law mass distribution for the heavier black hole in the BBHs as [40, 41, 42, 43]
+    $\begin{split}&P_{\rm BBH}(m_{\rm s})=Z_{\rm m}m_{\rm s}^{-\alpha}\times\\ &\mathcal{H}(m_{\rm s}-m_{\min})\mathcal{H}(m_{\max}-m_{\rm s}),\end{split}$ (13)
+    where $\mathcal{H}$ is the Heaviside step function and $Z_{\rm m}$ is the normalization constant as
+    $Z_{\rm m}=\left\{\begin{aligned} \frac{1-\alpha}{M_{\max}^{1-\alpha}-M_{\min}^{1-\alpha}},\quad\alpha\neq 1,\\ \frac{1}{\ln(M_{\max}/M_{\min})},\quad\alpha=1.\end{aligned}\right$. (14)
+    The heavier black hole $m_{1}$ satisfies the mass distribution as Eq.(13) with $\alpha=2.35$ in the range of $m_{1}\in[5,100]\penalty 10000\ M_{\odot}$, and the mass of $m_{2}$ uniformly distributes in the interval $m_{1}/18,m_{1}]$.
+For other parameters in GW sources, such as the spin ($s_{1}=s_{2}=0)$ and eccentricity ($e=0$), we adopt a fixed value.
+In addition, we take orbital inclination ($\iota\in[0,\pi]$), the coalescence phase ($\phi_{\rm c}\in[0,2\pi]$), location angles ($\theta\in[0,2\pi]$, $\phi\in[-\pi/2,\pi/2]$), and the polarisation angle ($\psi\in[0,\pi]$) as uniform distribution in GW waveform.
+We employ Python module PyCBC to produce IMRPhenomD waveform of frequency-domain for unlensed GW signals $\bar{h}_{0}(f)$ [57].
+    Based on the redshift distribution and mass distribution in above subsection, we can simulate $10^{4}$ detectable unlensed BBH events with $\rm{SNR}\geq 8$ for ET.
+The SNR of the unlensed GW signal can be computed as
+    ${\rm{SNR}_{0}}=\left\langle\bar{h}_{0}|\bar{h}_{0}\right\rangle=\sqrt{4\int\frac{|\bar{h}_{0}(f)|^{2}}{S_{\rm n}(f)}df},$ (15)
+    where $S_{\rm n}(f)$ is the power spectral density for ET.
+  • Generating the micro-lensing GW events: By assuming that the compact DM is distributed uniformly in comoving volume, the probability that a GW source at $z_{\rm s}$ is lensed by an intervening compact DM object is given by optical depth
+    $P_{\rm L}(z_{\rm s})=1-\exp(-\tau(z_{\rm s})).$ (16)
+    We identify a BBH GW signal as a lensed event when $P_{\rm L}(z_{\rm s})$ is larger than a random number uniformly distribution between 0 and 1.
+Therein the optical depth $\tau(z_{\rm s})$ can be written as
+    $\begin{split}\tau(z_{\rm s})=&\int_{0}^{+\infty}dm_{\rm l}\int_{0}^{z_{\rm s}}dz_{\rm l}\int_{0}^{y_{0}}dy\\ &\tau(m_{\rm l},z_{\rm l},y),\end{split}$ (17)
+    where $y_{0}$ is the maximum impact parameter.
+$\tau(m_{\rm l},z_{\rm l},y)$ is the differential optical depth evaluated at the mass of lens $m_{\rm l}$, redshift of lens $z_{\rm l}$ and impact parameter $y$, given by
+    $\begin{split}\tau(m_{\rm l},z_{\rm l},y)=&3f_{\rm CO}\Omega_{\rm DM}y\psi(m_{\rm l},\bm{p}_{\rm mf})\times\\ &\frac{H_{0}^{2}(1+z_{\rm l})^{2}}{H(z_{\rm l})}\frac{D_{\rm l}D_{\rm ls}}{D_{\rm s}},\end{split}$ (18)
+    where we set $f_{\rm CO}=10^{-2.5}$.
+  • Generating the parameters of lens: When a GW event is identified as a lensed one, the value of $z_{\rm l}$ is adopted from a probability distribution $P(z_{\rm l})$ as given in Eq. (18)
+    $P(z_{\rm l})=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{z_{\rm s}}dz_{\rm l}\penalty 10000\ \tau(m_{\rm l},z_{\rm l},y)}.$ (19)
+    In addition, the distribution $P(y)$ also can be obtained from Eq. (18) in the range of $y\in[0,y_{0}]$
+    $P(y)=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{y_{0}}dy\penalty 10000\ \tau(m_{\rm l},z_{\rm l},y)}=\frac{2}{y_{0}^{2}}y.$ (20)
+    We set $y_{0}=5$ as mentioned above.
+For extended lens mass distribution $\psi(m_{\rm l},\bm{p}_{\rm mf})$ which is the same as that in Eq.
+(28), we assume the maximum value of the extended power-law mass function is $M_{\max}=10^{3}\penalty 10000\ M_{\odot}$ and the minimum value is $M_{\min}=10\penalty 10000\ M_{\odot}$.
+The mass of the lens satisfies the power-law distribution described as
+    $\begin{split}\psi(m_{\rm l},\gamma)=&\mathcal{N}_{\rm pl}m_{\rm l}^{\gamma-1}\mathcal{H}(m_{\rm l}-M_{\min})\times\\ &\mathcal{H}(M_{\max}-m_{\rm l}),\end{split}$ (21)
+    where $\mathcal{N}_{\rm pl}$ is the normalization, and we set the power-law index $\gamma=1$.
+  • Computing lensed criterion: After generating unlensed and lensed GW waveform with simulated parameters for sources and lens, we use the mismatch $\epsilon$ and $\rm{SNR_{L}}$ as criteria for the selection of lensing GW following [36]
+    $\epsilon\times{\rm{SNR_{L}}}^{2}>1,$ (22)
+    where the mismatch $\epsilon$ is expressed as
+    $\epsilon=1-\frac{\left\langle\bar{h}_{\rm L}|\bar{h}_{0}\right\rangle}{\sqrt{\left\langle\bar{h}_{\rm L}|\bar{h}_{\rm L}\right\rangle\left\langle\bar{h}_{0}|\bar{h}_{0}\right\rangle}}.$ (23)
+  • Extract parameters of micro-lensing GWs: We use the Bayes’ theorem to forecast the constraints of the parameters $m_{1},m_{2},d_{\rm L}(z_{\rm s}),M_{\rm l}^{z},y]$ from lensed GW signals.
+The likelihood for each GW signals can be expressed as an inner product [51]
+    $p(d|\theta)\propto\exp\bigg[-\frac{1}{2}\left\langle d-\bar{h}(\theta)|d-\bar{h}(\theta)\right\rangle^{2}\bigg],$ (24)
+    where $\bar{h}(\theta)$ represents the model of micro-lensing GW waveform.
+Figure 2: Left: The proportion of detectable micro-lensing GW events for Eq.
+(34).
+The blue data points represent simulated values, while the solid red line indicates the corresponding polynomial fit to these data.
+Right: Solid lines with points represent inferred lens masses of 8 micro-lensing GWs at $95\%$ confidence levels with median value derived from the posterior distribution of redshifted lens mass $p(M_{\rm l}^{z}|d_{i})$ and luminosity-distance distribution of GW sources $p(d_{\rm L}(z_{\rm s})|d_{i})$ in Figure 1.
+The star points represent the true values of the simulated lens masses.
+From $10^{4}$ simulated BBH GW events, we identify 8 detectable micro-lensing GW signals.
+Moreover, we present the posterior distribution of the parameters of these events in Figure 1.
+As shown in the right panel of Figure 1, we compare the true simulated values of the luminosity distance for GW source and the redshift mass of lens with their posterior-derived counterparts.
+These results demonstrate good agreement within $2\sigma$ confidence levels.
+IV Hierarchical Bayesian Inference
+For population hyperparameters of compact DM $\Phi$ and $N_{\rm obs}$ detections of micro-lensing BBHs events $d=[d_{1},...,d_{N_{\rm obs}}]$ ($N_{\rm obs}=8$ from $10^{4}$ simulated GW events as the simulation result in Section III), the likelihood follows an in-homogeneous Poisson distribution without any measurement uncertainty and selection effect is
 $p(d|\Phi)\propto N(\Phi)^{N_{\rm obs}}e^{-N(\Phi)}\prod_{i}^{N_{\rm obs}}p_{\rm pop}(d_{i}|\Phi).$ (25)
@@
 In Eq.
-(9), $\sigma(m_{\rm l},z_{\rm l},z_{\rm s})$ is the lensing cross section
+(28), population hyperparameters of compact DM are defined as
+$\Phi\equiv[\bm{p}_{\rm mf},f_{\rm CO}].$ (29)
+and $\sigma(m_{\rm l},z_{\rm l},z_{\rm s})$ in Eq. (27) is the lensing cross section
 $\sigma(m_{\rm l},z_{\rm l},z_{\rm s})=\frac{4\pi m_{\rm l}D_{\rm l}D_{\rm ls}}{D_{\rm s}}y_{0}^{2}.$ (30)
-It should be emphasized that we take the impact parameter as $y_{0}=5$, because $y_{0}>5$ are difficult to be identified as micro-lensing signals [29, 31].
+It should be emphasized that we take the impact parameter as $y_{0}=5$, because when $y_{0}>5$, it’s difficult to identify the micro-lensing signal [41, 43].
 In addition, $p_{\rm pop}(m_{\rm l}|\Phi)$ is the normalized distribution of lens masses and written as
@@
 Meanwhile, $N_{\rm det}(\Phi)$ is the number of detectable micro-lensing GW events and can be defined as
-$\begin{split}N_{\rm det}(\Phi)=&\int dm_{\rm l}\int dz_{\rm s}\int_{0}^{z_{\rm s}}dz_{\rm l}\frac{dn(m_{\rm l},\Phi)}{dm_{\rm l}}\times\\ &\frac{d\chi(z_{\rm l})}{dz_{\rm l}}(1+z_{\rm l})^{2}\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})N_{\rm s}P_{\rm s}(z_{\rm s}),\end{split}$ (13)
-where $\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})$ is the cross section that micro-lensing signal can be detected and depends on Eq. (30) via
-$\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})=\frac{4\pi m_{\rm l}D_{\rm l}D_{\rm ls}}{D_{\rm s}}y_{\max}(\lambda)^{2}.$ (14)
+$\begin{split}N_{\rm det}(\Phi)=&\int dm_{\rm l}\int dz_{\rm s}\int_{0}^{z_{\rm s}}dz_{\rm l}\frac{dn(m_{\rm l},\Phi)}{dm_{\rm l}}\times\\ &\frac{d\chi(z_{\rm l})}{dz_{\rm l}}(1+z_{\rm l})^{2}\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})\times\\ &N_{\rm s}P_{\rm s}(z_{\rm s}),\end{split}$ (32)
+where $\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})$ is the cross section of the micro-lensing signal that can be detected depending on Eq.
+(22).
+The cross section $\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})$ is given by
+$\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})=\frac{4\pi m_{\rm l}D_{\rm l}D_{\rm ls}}{D_{\rm s}}y_{\max}(\lambda)^{2},$ (33)
 where $\lambda\in[M_{\rm l}^{z},z_{\rm s},m_{1},m_{2},$......] contains the parameter sets of the GW source and lens. Meanwhile, we can define the detected fraction $\xi(\Phi)$ of micro-lensing GWs as
@@
 In addition, $Z_{\mathcal{M}}$ is both the normalized factor and the Bayesian evidence for the population model $\mathcal{M}$.
-This normalized factor can be calculated from the integral of the numerator of Eq.
-(16) over $\Phi$, i.e.
-$Z_{\mathcal{M}}=\int d\Phi\leavevmode\nobreak\ p(d|\Phi)p(\Phi).$ (17)
-Figure 1: Left: The posterior distribution for one of lensing GW events in 8 detectable micro-lensing GW events.
-Right: The cross represents the range of values within $2\sigma$ confidence level of luminosity distance for GW sources and redshift mass for lens extracted from the posterior distribution of 8 detectable micro-lensing GW events.
-And the scattered points represent the true values of these micro-lensing events.
-III Simulation
-In this Section, we introduce the simulation of GWs and how to generate micro-lensing GW events.
-III.1 Simulation for model of GW sources
-Although understanding the origin of GWs of BBHs is still an open issue, we select BBH merger events as the GW sources for this investigation due to some reasons: These systems exhibit high signal-to-noise ratio (SNR), robust waveform templates, and high event rates.
-The redshift distribution of GW sources is drawn from the phenomenological fit $R_{\rm BBH}$ to the population synthesis rate
-$P_{\rm BBH}(z_{\rm s})=\frac{1}{Z_{\rm BBH}}\frac{R_{\rm BBH}(z_{\rm s})}{1+z_{\rm s}}\frac{dV_{\rm c}}{dz_{\rm s}},$ (18)
-where $Z_{\rm BBH}$ is the normalization constant, and $R_{\rm BBH}$ is the merger rate per unit of volume in the source frame and takes the form [30, 47]
-$R_{\rm BBH}(z_{\rm s})=R_{0}\int_{t_{\min}}^{t_{\max}}R_{\rm f}(t(z_{\rm s})-t_{\rm d})P(t_{\rm d})dt_{\rm d},$ (19)
-where $t_{\min}=50\leavevmode\nobreak\ {\rm Myr}$, $t_{\max}$ is set to the Hubble time, $P(t_{\rm d})\propto t_{\rm d}^{-1}$ is the time delay distribution which is set the power-law, and $R_{\rm f}(t(z_{\rm s})-t_{\rm d})$ is the Star Formation Rate (SFR) as [48, 30]
-$R_{\rm f}(z_{\rm s})=[1+(1+z_{\rm p})^{-\lambda-\kappa}]\frac{(1+z_{\rm s})^{\lambda}}{1+\bigg{(}\frac{1+z_{\rm s}}{1+z_{\rm p}}\bigg{)}^{\lambda+\kappa}},$ (20)
-where we set $\lambda=2.6$, $\kappa=3.6$, and $z_{\rm p}=2.2$. We use a power-law mass distribution for the heavier black hole in the BBHs as [28, 29, 30, 31]
-$P_{\rm BBH}(m_{\rm s})=Z_{\rm m}m_{\rm s}^{-\alpha}\mathcal{H}(m_{\rm s}-m_{\min})\mathcal{H}(m_{\max}-m_{\rm s}),$ (21)
-where $\mathcal{H}$ is the Heaviside step function and $Z_{\rm m}$ is the normalization constant as
-$Z_{\rm m}=\left\{\begin{aligned} \frac{1-\alpha}{M_{\max}^{1-\alpha}-M_{\min}^{1-\alpha}},\quad\alpha\neq 1,\\ \frac{1}{\ln(M_{\max}/M_{\min})},\quad\alpha=1.\end{aligned}\right$. (22)
-The heavier black hole $m_{1}$ satisfies the mass distribution as Eq.(21) with $\alpha=2.35$ in the range of $m_{1}\in[5,100]\leavevmode\nobreak\ M_{\odot}$, and the mass of $m_{2}$ uniformly distributes in the interval $m_{1}/18,m_{1}]$  [^1] For other parameters in GW sources, such as the orbital inclination and the merging orbital phase, we adopt a uniform distribution.
-In the following discussion, we do not show them, but we all take their influence into consideration..
-III.2 Simulation for micro-lensing events
-To ensure the physical rationality and computational efficiency of micro-lensing GW events simulation, we have developed a five-step process:
-  • Generating the detectable BBH events: Based on the redshift distribution and mass distribution in above subsection, we can simulate $10^{4}$ detectable unlensed BBH events with $\rm{SNR}\geq 8$ for ET.
-The SNR of the unlensed GW signal can be computed as
-    ${\rm{SNR}_{0}}=\left\langle\bar{h}_{0}|\bar{h}_{0}\right\rangle=\sqrt{4\int\frac{|\bar{h}_{0}(f)|^{2}}{S_{\rm n}(f)}df},$ (23)
-    where $S_{\rm n}(f)$ is the power spectral density for ET.
-  • Generating the micro-lensing GW events: By assuming that the compact DM is distributed uniformly in comving volume, the probability that a GW source at $z_{\rm s}$ is lensed by an intervening compact DM object can be given by optical depth as
-    $P_{\rm L}(z_{\rm s})=1-\exp(-\tau(z_{\rm s})).$ (24)
-    We identify a BBH GW signal as a lensed event when $P_{\rm L}(z_{\rm s})$ is larger than a random number uniformly distribution between 0 and 1.
-Therein the optical depth $\tau(z_{\rm s})$ can be written as
-    $\begin{split}\tau(z_{\rm s})=&\int_{0}^{+\infty}dm_{\rm l}\int_{0}^{z_{\rm s}}dz_{\rm l}\int_{0}^{y_{0}}dy\\ &\tau(m_{\rm l},z_{\rm l},y),\end{split}$ (25)
-    where $y_{0}$ is the maximum impact parameter.
-$\tau(m_{\rm l},z_{\rm l},y)$ is the differential optical depth for the mass of lens $m_{\rm l}$, redshift of lens $z_{\rm l}$ and impact parameter $y$, given by
-    $\begin{split}\tau(m_{\rm l},z_{\rm l},y)=&3f_{\rm CO}\Omega_{\rm DM}y\psi(m_{\rm l},\bm{p}_{\rm mf})\times\\ &\frac{H_{0}^{2}(1+z_{\rm l})^{2}}{H(z_{\rm l})}\frac{D_{\rm l}D_{\rm ls}}{D_{\rm s}},\end{split}$ (26)
-    where we set $f_{\rm CO}=10^{-2.5}$.
-  • Generating the parameters of lens: When a GW event is identified as lensed one, the value of $z_{\rm l}$ is adopted from a probability distribution $P(z_{\rm l})$ given by Eq. (26)
-    $P(z_{\rm l})=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{z_{\rm s}}dz_{\rm l}\leavevmode\nobreak\ \tau(m_{\rm l},z_{\rm l},y)}.$ (27)
-    In addition, the distribution $P(y)$ also can be obtained from Eq. (26) in the range of $y\in[0,y_{0}]$
-    $P(y)=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{y_{0}}dy\leavevmode\nobreak\ \tau(m_{\rm l},z_{\rm l},y)}=\frac{2}{y_{0}^{2}}y.$ (28)
-    We set $y_{0}=5$ as mentioned above.
-For extended mass distribution of lens mass $\psi(m_{\rm l},\bm{p}_{\rm mf})$ same as in Eq.
-(10), we assume the maximum value of the extended power-law mass function is $M_{\max}=10^{3}\leavevmode\nobreak\ M_{\odot}$ and the minimum value is $M_{\min}=10\leavevmode\nobreak\ M_{\odot}$.
-The mass of the lens satisfies the power-law distribution described as
-    $\begin{split}\psi(m_{\rm l},\gamma)=&\mathcal{N}_{\rm pl}m_{\rm l}^{\gamma-1}\mathcal{H}(m_{\rm l}-M_{\min})\times\\ &\mathcal{H}(M_{\max}-m_{\rm l}),\end{split}$ (29)
-    where $\mathcal{N}_{\rm pl}$ is the normalization, and we set the power-law index $\gamma=1$.
-  • Computing lensed criterion: After generating unlensed and lensed GW waveform with simulated parameters for sources and lens, we use the mismatch $\epsilon$ and $\rm{SNR_{L}}$ as criterion for selection of lensing GW following [25]
-    $\epsilon\times{\rm{SNR_{L}}}^{2}>1,$ (30)
-    where mismatch $\epsilon$ is expressed as
-    $\epsilon=1-\frac{\left\langle\bar{h}_{\rm L}|\bar{h}_{0}\right\rangle}{\sqrt{\left\langle\bar{h}_{\rm L}|\bar{h}_{\rm L}\right\rangle\left\langle\bar{h}_{0}|\bar{h}_{0}\right\rangle}}.$ (31)
-  • Extract parameters of micro-lensing GWs: We use the Bayes’ theorem to forecast the constraints of the parameters $m_{1},m_{2},d_{\rm L}(z_{\rm s}),M_{\rm l}^{z},y]$ from lensed GW signals.
-The likelihood for each GW signals can be written in the form of an inner product [39]
-    $p(d|\theta)\propto\exp\bigg{[}-\frac{1}{2}\left\langle d-\bar{h}(\theta)|d-\bar{h}(\theta)\right\rangle^{2}\bigg{]},$ (32)
-    where $\bar{h}(\theta)$ represents the model of micro-lensing GW waveform.
-Figure 2: Left: The proportion of detectable micro-lensing GW events for Eq.
-(15).
-The blue data points represent simulated values, while the solid red line indicates the corresponding polynomial fit to these data.
-Right: Solid lines with points represent inferred lens masses of 8 micro-lensing GWs at $95\%$ confidence levels with median value derived from the posterior distribution of redshifted lens mass $p(M_{\rm l}^{z}|d_{i})$ and luminosity-distance distribution of GW sources $p(d_{\rm L}(z_{\rm s})|d_{i})$ in Figure 1.
-The star points represent the true values of the simulated lens masses.
-From $10^{4}$ simulated GW events of BBHs, we identify 8 detectable micro-lensing GW signals.
-Moreover, we present the posterior distribution of the parameters of these events in Figure 1.
-As shown in the right panel of Figure 1, we compare the true simulated values of the luminosity distance for GW source and the redshift mass of lens with their posterior-derived counterparts.
-These results demonstrate good agreement within $2\sigma$ confidence levels.
-IV Results
+V Results
 The selection effects of gravitational lensing systems and the uncertainties in observed lens masses are presented in Figure 2.
@@
 The blue solid point represent the truth values for hyperparameters $\gamma=1,\log_{10}(f_{\rm CO})=-2.5]$ in our simulation.
-For the population hyperparameters of compact DM, we fix the minimum mass $M_{\min}=10\leavevmode\nobreak\ M_{\odot}$ and maximum mass $M_{\max}=10^{3}\leavevmode\nobreak\ M_{\odot}$ in the mass function of compact DM, and we adopt uniform priors for the power-law index $\gamma\in\mathcal{U}[-1,3]$ and abundance of compact DM $f_{\rm CO}\in\mathcal{U}[-4,-1]$ in Eq.
+For the population hyperparameters of compact DM, we fix the minimum mass $M_{\min}=10\penalty 10000\ M_{\odot}$ and maximum mass $M_{\max}=10^{3}\penalty 10000\ M_{\odot}$ in the mass function of compact DM, and we adopt uniform priors for the power-law index $\gamma\in\mathcal{U}[-1,3]$ and abundance of compact DM $f_{\rm CO}\in\mathcal{U}[-4,-1]$ in Eq.
 (35), respectively.
-Then we incorporate the above-mentioned 8 inferred lens masses from micro-lensing GW events into the EMCEE [49] with the posterior Eq.
-(16) to estimate population hyperparameters $\Phi\in[\gamma,\leavevmode\nobreak\ f_{\rm CO}]$.
+Then we incorporate the above-mentioned 8 inferred lens masses from micro-lensing GW events into the EMCEE [61] with the posterior in Eq.
+(35) to estimate population hyperparameters $\Phi\in[\gamma,\penalty 10000\ f_{\rm CO}]$.
 Our results are shown in Figure.
 3.
-We obtain that the best-fit values and $68\%$ confidence levels for the hyperparameters $\gamma,\leavevmode\nobreak\ f_{\rm CO}]$ are $\gamma=0.41^{+0.42}_{-0.37}$, $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$, respectively.
-As shown by the blue data point in Figure 3, which indicate the truth values for hyperparameter $\Phi\in[\gamma=1,\leavevmode\nobreak\ \log_{10}(f_{\rm CO})=-2.5]$ in our simulations, our posterior estimates demonstrate consistency with these input truth values within $2\sigma$ confidence levels.
-V Conclusion
+We obtain that the best-fit values and $68\%$ confidence levels for the hyperparameters $\gamma,\penalty 10000\ f_{\rm CO}]$ are $\gamma=0.41^{+0.42}_{-0.37}$, $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$, respectively.
+As shown by the blue data point in Figure 3, which indicate the truth values for hyperparameter $\Phi\in[\gamma=1,\penalty 10000\ \log_{10}(f_{\rm CO})=-2.5]$ in our simulations, our posterior estimates demonstrate consistency with these input truth values within $2\sigma$ confidence levels.
+VI Conclusion
 In this paper, we present a comprehensive methodology for: (1) simulating GW events mciro-lensed by compact DM, (2) extracting information about both the GW sources and lens for each micro-lensing events, and (3) constraining the population properties of compact DM from these events.
@@
 However, only 8 micro-lensing GW events were detected, which explains why the derived constraints on both $f_{\rm CO}$ and $\gamma$ are lower than the true input values.
-The implementation of our methodology will play a pivotal role in future observations of micro-lensing GWs: it enables not only the precise extraction of population characteristics of compact DM, but also provides a pathway to probe their fundamental nature—whether they stem from primordial black holes, exotic compact objects, or other non-particle dark matter candidates.
+The implementation of our methodology will play a pivotal role in future observations of micro-lensing GWs: it enables not only the precise extraction of population characteristics of compact DM, but also provides a pathway to probe their fundamental nature, whether originating from primordial black holes, exotic compact objects, or other non-particle dark matter candidates.
+Although our method has established a preliminary framework for constraining dark matter using future gravitational wave data, it still needs to account for additional factors to better approximate the real physical process, such as
+  • While our study focuses exclusively on the point-mass model, compact DM in the universe may exist in other extended halo structures, such as “dressed PBH” [62, 63, 64, 65, 66].
+If an observed signal is a true micro-lensing waveform, using an incorrect template model will produce a biased result.
+Therefore, following the detection of a micro-lensing GW event, it is crucial to accurately identify the specific lens model involved [35, 51].
+  • Conversely, the second case arises when an observed waveform is not lensed but is spuriously identified as a micro-lensing signal.
+A prime example is the analysis of GW from a binary in an elliptical orbit using templates designed for quasi-circular orbits.
+In such instances, the micro-lensing analysis may mistakenly interpret the elliptical orbit signature as evidence of lensing.
+Consequently, additional methods are required to break this degeneracy with the parameters of lens, for example, employing the Kramers-Kronig (KK) relation [67].
+  • To date, no GW event has shown unambiguous signs of micro-lensing.
+Future detections will require more complete waveform models to robustly interpret any deviations and to assess the micro-lensing contribution from known stellar populations [68, 69, 70, 71, 72].
+Ultimately, determining whether we can reliably identify dark matter substructures, distinct from the effects of known baryonic matter, remains an open challenge.
 The constraints on compact DM will improve significantly due to the rapidly growing number of micro-lensing GW detections from upcoming surveys and the enhanced horizon distance of next-generation GW detectors.
@@
 Such joint analyses are expected to play a pivotal role in unraveling the nature of compact DM and their formation mechanisms, shedding light on physics in the early universe.
-VI Acknowledgements
+VII Acknowledgements
 This work is supported by National Key R&D Program of China under Grant No.2024YFC2207400; China National Postdoctoral Program for Innovative Talents under Grant No.BX20230271; National Key R&D Program of China Grant No.